# MMEdge: Accelerating On-device Multimodal Inference via Pipelined Sensing and Encoding

### v2

@@ Abstract.
Real-time multimodal inference on resource-constrained edge devices is essential for applications such as autonomous driving, human-computer interaction, and mobile health. However, prior work often overlooks the tight coupling between sensing dynamics and model execution, as well as the complex inter-modality dependencies. In this paper, we propose MMEdge, a new on-device multimodal inference framework based on pipelined sensing and encoding. Instead of waiting for complete sensor inputs, MMEdge decomposes the entire inference process into a sequence of fine-grained sensing and encoding units, allowing computation to proceed incrementally as data arrive. MMEdge also introduces a lightweight but effective temporal aggregation module that captures rich temporal dynamics across different pipelined units to maintain accuracy performance. Such pipelined design also opens up opportunities for fine-grained cross-modal optimization and early decision-making during inference. To further enhance system performance under resource variability and input data complexity, MMEdge incorporates an adaptive multimodal configuration optimizer that dynamically selects optimal sensing and model configurations for each modality under latency constraints, and a cross-modal speculative skipping mechanism that bypasses future units of slower modalities when early predictions reach sufficient confidence. We evaluate MMEdge using two public multimodal datasets and deploy it on a real-world unmanned aerial vehicle (UAV)-based multimodal testbed. The results show that MMEdge significantly reduces end-to-end latency while maintaining high task accuracy across various system and data dynamics A video demonstration of MMEdge’s performance in real world is available at https://youtu.be/n36M9ho2z9o.  [^1] Code available at: https://github.com/HKUST-MINSys-Lab/MMEdge. Accepted by SenSys 2026.
copyright: none
@@ 1. Introduction
Multimodal sensing and learning systems are increasingly adopted in real-time applications such as autonomous driving (Xie et al., 2023; Wu et al., 2024), fall detection (Chen et al., 2020), and wearable interaction (Mayberry et al., 2015; Pang et al., 2023). Such systems continuously collect, process and fuse data from heterogeneous sensors (e.g., cameras, microphones, IMUs) to provide improved performance for complex tasks in dynamic environments. For example, RCBEVDet (Lin et al., 2024) shows that integrating radar with camera can significantly enhance object detection accuracy in autonomous driving. While offloading computation to the cloud has been widely adopted in distributed multimodal sensing systems, it raises growing concerns about user’s data privacy. Moreover, such systems often experience significant latency due to unpredictable network dynamics (Li et al., 2021; Huang et al., 2023) and the high bandwidth demands of data-intensive sensors like LiDAR (Wu et al., 2024). As a result, there is an increasing demand for efficient on-device multimodal systems, which enables end-to-end processing of sensor data locally without transmitting sensor data.
@@
However, deploying end-to-end multimodal perception systems on devices presents several major challenges, particularly under limited and dynamic resource conditions. First, there is an inherent dependency between the sensing and model inference phases when they share computational resources on the same device, making isolated optimization strategies ineffective. For instance, increasing model complexity or enhancing sensing granularity can both improve prediction accuracy, but also introduce higher latency, necessitating careful coordination between the two. Second, multimodal systems also exhibit strong inter-modality dependencies, where both fusion accuracy and latency depend on the joint optimization of all modalities. For example, Shuai et al. (Shuai et al., 2021) show that human detection accuracy in radar-camera systems depends on synchronized multimodal inputs, especially under varying lighting conditions, rather than relying on any single modality alone. Moreover, there will be asynchronous processing delays across modalities, which often leads to idle waiting time and inefficient resource utilization. In such cases, faster modalities are forced to wait for slower ones, increasing end-to-end latency. These challenges motivate a unified system design that dynamically coordinates the sensing and inference configurations across different modalities at run time to optimize both accuracy and efficiency.
@@
Unfortunately, most prior work on on-device multimodal systems has focused on optimizing either the sensing or the model inference stage in isolation, without considering their interdependencies. Some methods focus on mitigating delays caused by asynchronous sensor inputs. For example, (Li et al., 2021) introduced modality imputation to synthesize missing data from slow modalities using fast ones to reduce waiting time, while modality gating methods (Hou et al., 2023; Mayberry et al., 2015) selectively skip uninformative sensor modalities before inference to improve efficiency. On the other hand, several methods aim at optimizing the model inference process itself through early exit strategies (Cao et al., 2022) or expert model selection (Liu et al., 2024). However, our results indicate that optimizing sensing or inference in isolation performs poorly under real-world system and data dynamics. Although some distributed systems consider sensing-inference interactions (Li et al., 2021; Wu et al., 2024), they are designed to mitigate the impact of network dynamics during sensor data transmission (Rastikerdar et al., 2024; Huang et al., 2023), which cannot address the shared computational resource constraints between sensing and inference that are intrinsic to on-device multimodal systems.
@@ 2. Related Work
Multimodal sensing and learning systems. Multimodal sensing and learning systems are becoming increasingly prevalent in real-world applications. These systems continuously collect and fuse data from different sensor modalities to enhance performance in complex, dynamic environments. For instance, ADMarker (Ouyang et al., 2024) fuses depth, radar, and audio data to detect daily activities for monitoring digital biomarkers of Alzheimer’s Disease. Similarly, autonomous driving (Shi et al., 2022; He et al., 2023) and robotic (Dellaert and Contributors, 2022; Kümmerle et al., 2011; Wang et al., 2023a) systems leverage multiple sensor modalities to expand sensing coverage and maintain robust performance under varying environmental conditions. However, most work in this space is focused on improving the accuracy of multimodal fusion, often overlooking optimizing inference latency which a critical factor for time-sensitive applications in practical deployments.
@@
Accelerating on-device inference. Reducing on-device inference latency is crucial for deploying deep models on resource-constrained devices. Prior works explore dynamic adaptation strategies, including early-exit mechanisms (Leontiadis et al., 2021; Jeon et al., 2023), model or input reconfiguration (Rastikerdar et al., 2024; Lee et al., 2024; Mayberry et al., 2015; Yang et al., 2020; Lu et al., 2010; Naderiparizi et al., 2017), and parallel execution frameworks (Wei et al., 2023; Ling et al., 2022). For instance, CACTUS (Rastikerdar et al., 2024) selects micro-classifiers based on context, while Glimpse (Naderiparizi et al., 2017) prunes irrelevant inputs using auxiliary sensors. However, these methods focus on single-modality inference and neglect the end-to-end coupling between sensing and computation. In contrast, MMEdge introduces pipelined sensing and encoding for multimodal systems, bridging sensor data acquisition and model execution in a latency-aware manner.
@@
Efficient multimodal inference. Achieving real-time multimodal sensing and inference on devices is challenging due to the overhead of processing multiple modalities. Some works (Mayberry et al., 2015; Cao et al., 2022; Xie et al., 2023) leverage prior knowledge from one modality to guide inference in another. For example, Glimpse (Naderiparizi et al., 2017) uses auxiliary sensors to filter out irrelevant modalities, while CIDER (Mayberry et al., 2015) employs infrared sensors to dynamically switch vision model configurations. Other approaches aim to achieve adaptive multimodal fusion during inference. MobiVQA (Cao et al., 2022) adjusts visual processing branches based on textual information, and SMG (Hou et al., 2023) introduces a lightweight fusion method that selects informative features from both modalities for efficient inference. However, most of these works focus on optimizing either the sensing or inference stage in isolation. Although some distributed systems consider sensing-inference interactions (Li et al., 2021; Wu et al., 2024), they are designed to mitigate the impact of network dynamics during sensor data transmission (Huang et al., 2023; Rastikerdar et al., 2024), which cannot address the shared computational resource constraints between sensing and inference that are intrinsic to on-device multimodal systems.
@@ 3.1. Understanding On-Device Multimodal Systems
Achieving real-time multimodal sensing and inference on devices is important for applications that need reside data locally due to privacy concerns and require timely predictions. These systems typically perform end-to-end processing of synchronized multimodal streaming data within a time window to provide accurate perception and decision-making. For example, autonomous vehicles continuously collect data from cameras and mmWave radar, and must fuse this synchronized multimodal data for perception (Xie et al., 2023) and planning (Hu et al., 2023), often within 100 ms to enable timely control decisions (Luo, 2019). Another example is interactive systems that rely on audio-visual streams for tasks such as speech recognition (Petridis et al., 2018) and gesture recognition (Shi et al., 2021). Enabling low-latency multimodal sensing and inference directly on devices is essential for delivering responsive and preserving user data privacy.
@@
Traditional end-to-end inference pipeline. Figure 1 illustrates a typical on-device multimodal inference system, comprising the following stages: (1) Data Collection and Preprocessing. Sensors (e.g., cameras, microphones) capture raw physical signals. These signals are then preprocessed through operations like normalization, rescaling, and cropping to ensure compatibility with downstream deep learning models. (2) Unimodal Encoding. Preprocessed data from each modality is passed through different unimodal encoders (e.g., ResNet (He et al., 2016), Transformer (Vaswani, 2017)) to extract compact and informative unimodal feature embeddings. (3) Multimodal Fusion. The unimodal embeddings are fused via concatenation, attention mechanisms, or transformer-based fusion, to form a unified multimodal representation that aligns features across modalities. (4) Prediction. The fused representation is then used for task-specific inference, such as classification, detection, or generation, typically through MLP layers or decoders in multimodal models.
Challenges. We conduct a motivational study to examine the challenges of optimizing inference latency in on-device multimodal systems. Specifically, we evaluate an audio-visual speech recognition task using the Lip Reading in the Wild dataset (Chung and Zisserman, 2017). The experiments are conducted on an NVIDIA Jetson Xavier NX (16GB memory), configured in 2-core, 10W power mode to simulate a low-resource deployment environment. We use ResNet-50 for video backbone and a 2-layer CNN for audio encoder, and an attention-based module for multimodal fusion. Each sample consists of a 1-second recording, comprising video at 30 FPS, and audio at 16 kHz segmented into chunks of 800 samples. We simulate data collection by loading data at fixed intervals matching the original sensor frame rate. A background thread continuously collects data to emulate realistic sensing overheads. Figure 2 shows the measured latency across different stages of the end-to-end inference process.
@@ 3.2. Potential of Pipelined Sensing and Encoding
Comparison between the traditional and pipelined framework. We implement the pipelined inference framework using the same experimental setup as the traditional framework described in Section 3.1 on the audio-visual speech recognition task. For a fair comparison, both frameworks use the same backbone models. In the traditional setup, we use ResNet-50 (He et al., 2016) for video encoding and a 2-layer CNN for audio encoding. In the pipelined framework, since each unit contains less temporal context (e.g., only one frame for video modality), we adopt a 2D ResNet without temporal modeling to better suit the finer granularity of input.
@@ 5.1.2. Efficient Temporal Aggregation
Alternating Temporal Shift. In traditional sequential inference pipelines, feature encoders have access to the entire input sequence, allowing them to capture global temporal dependencies during feature extraction. In contrast, our pipelined design processes each data unit (e.g., a video frame or audio chunk) independently upon arrival. While this approach reduces latency, it will also limits the temporal context available to each unit. To address this limitation, we introduce an alternating temporal shift module that enables each unit to incorporate contextual information from its past and future neighbors. Inspired by temporal shift mechanisms (Lin et al., 2019), our method alternates feature channels along the temporal dimension, facilitating lightweight context propagation across adjacent units.
@@ 5.2.2. Problem Formulation
In the objective formulation of the problem (i.e., Equation 4), $A(\cdot)$ denotes an accuracy predictor that estimates the expected performance of different configurations on varying data inputs. However, directly performing this estimation from raw multimodal sensor data $x$ would require a large and computationally expensive model, making it impractical for real-time optimization in edge deployment. To address this, we propose a lightweight accuracy predictor that operates on informative features extracted from the raw multimodal data. Specifically, we introduce two auxiliary metrics designed to capture consistent and complementary information across modalities, which are important to influence the effectiveness of multimodal learning systems (Ouyang et al., 2022; Wu et al., 2024). These metrics enable the accuracy predictor to make accurate performance estimates without incurring significant computational overhead. Specifically, we define the two types of multimodal information as follows:
@@ 6.1. Testbed and Datasets
Real-world UAV Testbed. We built a UAV-based multimodal testbed and deployed MMEdge on it to evaluate end-to-end performance under resource constraints in real-world, time-critical applications. As shown in Figure 8, the testbed integrates an Intel RealSense D435iF camera (Int, [n. d.]) and a TI IWR1443BOOST mmWave radar (Texas Instruments, 2018) for multimodal data collection, both connected to an onboard NVIDIA Jetson Orin Nano edge computer (Corporation, 2024a). The task is to track human subjects on the ground using camera and radar data captured by the UAV in flight [^2] All the data collection was approved by the Institutional Review Board (IRB) of the authors’ institution.
@@
Public datasets. We also evaluate MMEdge on two public multimodal datasets: Lip Reading in the Wild (LRW) (Chung and Zisserman, 2017) and NuScenes-Mini-QA (Qian et al., 2024). LRW is a large-scale audio-visual dataset for speech recognition based on audio and video data, which includes over 500 spoken word classes and contains approximately 55,000 data samples. NuScenes-Mini-QA is a multimodal question answering (QA) dataset derived from the NuScenes autonomous driving scenes. It features temporal sequences of RGB images and LiDAR point clouds, as well as corresponding QA pairs, enabling question answering tasks related to dynamic driving scenarios.

### v5

@@ Abstract.
Real-time multimodal inference on resource-constrained edge devices is essential for applications such as autonomous driving, human-computer interaction, and mobile health. However, prior work often overlooks the tight coupling between sensing dynamics and model execution, as well as the complex inter-modality dependencies. In this paper, we propose MMEdge, a new on-device multimodal inference framework based on pipelined sensing and encoding. Instead of waiting for complete sensor inputs, MMEdge decomposes the entire inference process into a sequence of fine-grained sensing and encoding units, allowing computation to proceed incrementally as data arrive. MMEdge also introduces a lightweight but effective temporal aggregation module that captures rich temporal dynamics across different pipelined units to maintain accuracy performance. Such pipelined design also opens up opportunities for fine-grained cross-modal optimization and early decision-making during inference. To further enhance system performance under resource variability and input data complexity, MMEdge incorporates an adaptive multimodal configuration optimizer that dynamically selects optimal sensing and model configurations for each modality under latency constraints, and a cross-modal speculative skipping mechanism that bypasses future units of slower modalities when early predictions reach sufficient confidence. We evaluate MMEdge using two public multimodal datasets and deploy it on a real-world unmanned aerial vehicle (UAV)-based multimodal testbed. The results show that MMEdge significantly reduces end-to-end latency while maintaining high task accuracy across various system and data dynamics. A video demonstration of MMEdge’s performance in real world is available at https://youtu.be/n36M9ho2z9o.  [^1] Code available at: https://github.com/HKUST-MINSys-Lab/MMEdge. Accepted by SenSys 2026.
Efficient ML, Multimodal System, Edge AI
copyright: acmcopyrightjournalyear: 2026conference: The 24th ACM/IEEE International Conference on Embedded Artificial Intelligence and Sensing Systems; May 11–14, 2025; Saint-Malo, Francebooktitle: The 24th ACM/IEEE International Conference on Embedded Artificial Intelligence and Sensing Systems (ACM SenSys ’26), May 11–14, 2026, Saint-Malo, Francedoi: 10.1145/xxxxxxx.xxxxxisbn: nnn-n-nnnn-nnnn-n/nn/nnccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learningccs: Computer systems organization Embedded and cyber-physical systems
@@ 1. Introduction
Multimodal sensing and learning systems are increasingly adopted in real-time applications such as autonomous driving (xie2023timely; wu2024adaflow), fall detection (chen2020fall), and wearable interaction (mayberry2015cider; pang2023ubipose). Such systems continuously collect, process and fuse data from heterogeneous sensors (e.g., cameras, microphones, IMUs) to provide improved performance for complex tasks in dynamic environments. For example, RCBEVDet (lin2024rcbevdet) shows that integrating radar with camera can significantly enhance object detection accuracy in autonomous driving. While offloading computation to the cloud has been widely adopted in distributed multimodal sensing systems, it raises growing concerns about user’s data privacy. Moreover, such systems often experience significant latency due to unpredictable network dynamics (li2021low; huang2023re) and the high bandwidth demands of data-intensive sensors like LiDAR (wu2024adaflow). As a result, there is an increasing demand for efficient on-device multimodal systems, which enables end-to-end processing of sensor data locally without transmitting sensor data.
@@
However, deploying end-to-end multimodal perception systems on devices presents several major challenges, particularly under limited and dynamic resource conditions. First, there is an inherent dependency between the sensing and model inference phases when they share computational resources on the same device, making isolated optimization strategies ineffective. For instance, increasing model complexity or enhancing sensing granularity can both improve prediction accuracy, but also introduce higher latency, necessitating careful coordination between the two. Second, multimodal systems also exhibit strong inter-modality dependencies, where both fusion accuracy and latency depend on the joint optimization of all modalities. For example, Shuai et al. (shuai2021millieye) show that human detection accuracy in radar-camera systems depends on synchronized multimodal inputs, especially under varying lighting conditions, rather than relying on any single modality alone. Moreover, there will be asynchronous processing delays across modalities, which often leads to idle waiting time and inefficient resource utilization. In such cases, faster modalities are forced to wait for slower ones, increasing end-to-end latency. These challenges motivate a unified system design that dynamically coordinates the sensing and inference configurations across different modalities at run time to optimize both accuracy and efficiency.
@@
Unfortunately, most prior work on on-device multimodal systems has focused on optimizing either the sensing or the model inference stage in isolation, without considering their interdependencies. Some methods focus on mitigating delays caused by asynchronous sensor inputs. For example, (li2021low) introduced modality imputation to synthesize missing data from slow modalities using fast ones to reduce waiting time, while modality gating methods (hou2023smg; mayberry2015cider) selectively skip uninformative sensor modalities before inference to improve efficiency. On the other hand, several methods aim at optimizing the model inference process itself through early exit strategies (cao2022mobivqa) or expert model selection (liu2024deepseek). However, our results indicate that optimizing sensing or inference in isolation performs poorly under real-world system and data dynamics. Although some distributed systems consider sensing-inference interactions (li2021low; wu2024adaflow), they are designed to mitigate the impact of network dynamics during sensor data transmission (rastikerdar2024cactus; huang2023re), which cannot address the shared computational resource constraints between sensing and inference that are intrinsic to on-device multimodal systems.
@@ 2. Related Work
Multimodal sensing and learning systems. Multimodal sensing and learning systems are becoming increasingly prevalent in real-world applications. These systems continuously collect and fuse data from different sensor modalities to enhance performance in complex, dynamic environments. For instance, ADMarker (ouyang2024admarker) fuses depth, radar, and audio data to detect daily activities for monitoring digital biomarkers of Alzheimer’s Disease. Similarly, autonomous driving (shi2022vips; he2023vi) and robotic (gtsam; kummerle2011g; wang2023pypose) systems leverage multiple sensor modalities to expand sensing coverage and maintain robust performance under varying environmental conditions. However, most work in this space is focused on improving the accuracy of multimodal fusion, often overlooking optimizing inference latency which a critical factor for time-sensitive applications in practical deployments.
@@
Accelerating on-device inference. Reducing on-device inference latency is crucial for deploying deep models on resource-constrained devices. Prior works explore dynamic adaptation strategies, including early-exit mechanisms (leontiadis2021s; jeon2023harvnet), model or input reconfiguration (rastikerdar2024cactus; lee2024panopticus; mayberry2015cider; yang2020resolution; lu2010jigsaw; naderiparizi2017glimpse), and parallel execution frameworks (wei2023nn; ling2022blastnet). For instance, CACTUS (rastikerdar2024cactus) selects micro-classifiers based on context, while Glimpse (naderiparizi2017glimpse) prunes irrelevant inputs using auxiliary sensors. However, these methods focus on single-modality inference and neglect the end-to-end coupling between sensing and computation. In contrast, MMEdge introduces pipelined sensing and encoding for multimodal systems, bridging sensor data acquisition and model execution in a latency-aware manner.
@@
Efficient multimodal inference. Achieving real-time multimodal sensing and inference on devices is challenging due to the overhead of processing multiple modalities. Some works (mayberry2015cider; cao2022mobivqa; xie2023timely) leverage prior knowledge from one modality to guide inference in another. For example, Glimpse (naderiparizi2017glimpse) uses auxiliary sensors to filter out irrelevant modalities, while CIDER (mayberry2015cider) employs infrared sensors to dynamically switch vision model configurations. Other approaches aim to achieve adaptive multimodal fusion during inference. MobiVQA (cao2022mobivqa) adjusts visual processing branches based on textual information, and SMG (hou2023smg) introduces a lightweight fusion method that selects informative features from both modalities for efficient inference. However, most of these works focus on optimizing either the sensing or inference stage in isolation. Although some distributed systems consider sensing-inference interactions (li2021low; wu2024adaflow), they are designed to mitigate the impact of network dynamics during sensor data transmission (huang2023re; rastikerdar2024cactus), which cannot address the shared computational resource constraints between sensing and inference that are intrinsic to on-device multimodal systems.
@@ 3.1. Understanding On-Device Multimodal Systems
Achieving real-time multimodal sensing and inference on devices is important for applications that need reside data locally due to privacy concerns and require timely predictions. These systems typically perform end-to-end processing of synchronized multimodal streaming data within a time window to provide accurate perception and decision-making. For example, autonomous vehicles continuously collect data from cameras and mmWave radar, and must fuse this synchronized multimodal data for perception (xie2023timely) and planning (hu2023planning), often within 100 ms to enable timely control decisions (luo2019time). Another example is interactive systems that rely on audio-visual streams for tasks such as speech recognition (petridis2018end) and gesture recognition (shi2021face). Enabling low-latency multimodal sensing and inference directly on devices is essential for delivering responsive and preserving user data privacy.
@@
Traditional end-to-end inference pipeline. Figure 1 illustrates a typical on-device multimodal inference system, comprising the following stages: (1) Data Collection and Preprocessing. Sensors (e.g., cameras, microphones) capture raw physical signals. These signals are then preprocessed through operations like normalization, rescaling, and cropping to ensure compatibility with downstream deep learning models. (2) Unimodal Encoding. Preprocessed data from each modality is passed through different unimodal encoders (e.g., ResNet (he2016deep), Transformer (vaswani2017attention)) to extract compact and informative unimodal feature embeddings. (3) Multimodal Fusion. The unimodal embeddings are fused via concatenation, attention mechanisms, or transformer-based fusion, to form a unified multimodal representation that aligns features across modalities. (4) Prediction. The fused representation is then used for task-specific inference, such as classification, detection, or generation, typically through MLP layers or decoders in multimodal models.
Challenges. We conduct a motivational study to examine the challenges of optimizing inference latency in on-device multimodal systems. Specifically, we evaluate an audio-visual speech recognition task using the Lip Reading in the Wild dataset (chung2017lip). The experiments are conducted on an NVIDIA Jetson Xavier NX (16GB memory), configured in 2-core, 10W power mode to simulate a low-resource deployment environment. We use ResNet-50 for video backbone and a 2-layer CNN for audio encoder, and an attention-based module for multimodal fusion. Each sample consists of a 1-second recording, comprising video at 30 FPS, and audio at 16 kHz segmented into chunks of 800 samples. We simulate data collection by loading data at fixed intervals matching the original sensor frame rate. A background thread continuously collects data to emulate realistic sensing overheads. Figure 2 shows the measured latency across different stages of the end-to-end inference process.
@@ 3.2. Potential of Pipelined Sensing and Encoding
Comparison between the traditional and pipelined framework. We implement the pipelined inference framework using the same experimental setup as the traditional framework described in Section 3.1 on the audio-visual speech recognition task. For a fair comparison, both frameworks use the same backbone models. In the traditional setup, we use ResNet-50 (he2016deep) for video encoding and a 2-layer CNN for audio encoding. In the pipelined framework, since each unit contains less temporal context (e.g., only one frame for video modality), we adopt a 2D ResNet without temporal modeling to better suit the finer granularity of input.
@@ 5.1.2. Efficient Temporal Aggregation
Alternating Temporal Shift. In traditional sequential inference pipelines, feature encoders have access to the entire input sequence, allowing them to capture global temporal dependencies during feature extraction. In contrast, our pipelined design processes each data unit (e.g., a video frame or audio chunk) independently upon arrival. While this approach reduces latency, it will also limits the temporal context available to each unit. To address this limitation, we introduce an alternating temporal shift module that enables each unit to incorporate contextual information from its past and future neighbors. Inspired by temporal shift mechanisms (lin2019tsm), our method alternates feature channels along the temporal dimension, facilitating lightweight context propagation across adjacent units.
@@ 5.2.2. Problem Formulation
In the objective formulation of the problem (i.e., Equation 4), $A(\cdot)$ denotes an accuracy predictor that estimates the expected performance of different configurations on varying data inputs. However, directly performing this estimation from raw multimodal sensor data $x$ would require a large and computationally expensive model, making it impractical for real-time optimization in edge deployment. To address this, we propose a lightweight accuracy predictor that operates on informative features extracted from the raw multimodal data. Specifically, we introduce two auxiliary metrics designed to capture consistent and complementary information across modalities, which are important to influence the effectiveness of multimodal learning systems (ouyang2022cosmo; wu2024adaflow). These metrics enable the accuracy predictor to make accurate performance estimates without incurring significant computational overhead. Specifically, we define the two types of multimodal information as follows:
@@ 6.1. Testbed and Datasets
Real-world UAV Testbed. We built a UAV-based multimodal testbed and deployed MMEdge on it to evaluate end-to-end performance under resource constraints in real-world, time-critical applications. As shown in Figure 8, the testbed integrates an Intel RealSense D435iF camera (IntelRealSenseD435i) and a TI IWR1443BOOST mmWave radar (TI_IWR1443_2018) for multimodal data collection, both connected to an onboard NVIDIA Jetson Orin Nano edge computer (jetson_orin). The task is to track human subjects on the ground using camera and radar data captured by the UAV in flight [^2] All the data collection was approved by the Institutional Review Board (IRB) of the authors’ institution.
@@
Public datasets. We also evaluate MMEdge on two public multimodal datasets: Lip Reading in the Wild (LRW) (chung2017lip) and NuScenes-Mini-QA (qian2024nuscenes). LRW is a large-scale audio-visual dataset for speech recognition based on audio and video data, which includes over 500 spoken word classes and contains approximately 55,000 data samples. NuScenes-Mini-QA is a multimodal question answering (QA) dataset derived from the NuScenes autonomous driving scenes. It features temporal sequences of RGB images and LiDAR point clouds, as well as corresponding QA pairs, enabling question answering tasks related to dynamic driving scenarios.